# Efficient Prompting for LLM-based Generative Internet of Things

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capacities on various tasks, and integrating the capacities of LLMs into the Internet of Things (IoT) applications has drawn much research attention recently. Due to security concerns, many institutions avoid accessing state-of-the-art commercial LLM services, requiring the deployment and utilization of open-source LLMs in a local network setting. However, open-source LLMs usually have more limitations regarding their performance, such as their arithmetic calculation and reasoning capacities, and practical systems of applying LLMs to IoT have yet to be well-explored. Therefore, we propose a text-based generative IoT (GIoT) system deployed in the local network setting in this study. To alleviate the limitations of LLMs and provide service with competitive performance, we apply prompt engineering methods to enhance the capacities of the open-source LLMs, design a Prompt Management Module and a Post-processing Module to manage the tailored prompts for different tasks and process the results generated by the LLMs. To demonstrate the effectiveness of the proposed system, we discuss a challenging Table Question Answering (Table-QA) task as a case study of the proposed system, as tabular data is usually more challenging than plain text because of their complex structures, heterogeneous data types and sometimes huge sizes. We conduct comprehensive experiments on two popular Table-QA datasets, and the results show that our proposal can achieve competitive performance compared with state-of-the-art LLMs, demonstrating that the proposed LLM-based GIoT system can provide competitive performance with tailored prompting methods and is easily extensible to new tasks without training.
@@
Despite the remarkable capacities of LLMs, many issues need to be considered when integrating LLMs into IoT systems for an LLM-based GIoT system. First, following the Scaling Law [9], the capacities of an LLM are growing with its number of parameters and the scale of the training dataset, making LLMs often have a tremendous number of parameters, which lead to high hardware requirements for the training and inference. For example, popular open-source LLMs providing competitive performance on some public benchmarks often have around 70 billion parameters, such as Mixtral-8*7B [10] and Llama-3-70B [11]. These numbers of parameters make it not practical to deploy LLMs on edge IoT devices. Even though deploying LLMs on the edge and cloud servers can be an option, the efficiency of an LLM-based GIoT system still needs to be carefully considered. Besides the hardware requirements and efficiency issues, data privacy and security issues hinder many institutions from accessing commercial state-of-the-art LLMs. For example, for a healthcare IoT application collecting private data from patients, it is necessary to avoid uploading collected data to public, commercial LLM services for data analysis because of privacy issues. Therefore, a safe, transparent, and controllable open-source LLM deployed in a local network is more suitable for many institutions, even though commercial LLMs can provide better services. At last, the performance and the scalability of a LLM-based GIoT system are another two critical considerations because both commercial and open-source LLMs have some inherent limitations [12, 13, 14], such as their hallucination issues, limited reasoning capacities for complex tasks. Typically, prompting methods and fine-tuning are two directions that can alleviate these limitations. Specifically, prompting methods improve the performance of an LLM by designing tailored prompts for different tasks to elicit the capacities of an LLM. For example, Chain of Thought (CoT) [14] is a popular prompting method to improve the LLM’s reasoning performance by providing reasoning rationales. Program of Thoughts (PoT) [13] is another prompting method generating Python code to offload the reasoning and calculation tasks to the Python interpreter. Since these prompting methods focus on tailoring task-specific prompts for LLMs, they often lead to longer inference time because of their larger number of prompting tokens. By contrast, fine-tuning an LLM for a task can also improve the performance but requires label datasets and computation resources for the model training. As discussed in some studies [15, 16], fine-tuning an LLM for unseen tasks can increase the model’s bias and degrade its general capacities.
@@ II-B Table Question Answering
Since this study uses the Table-QA task as the case study of the proposed LLM-based GIoT system, we include recent studies using LLM to solve the Table-QA problem in this section. Specifically, most of the studies applying LLMs to the Table-QA task can be categorized into instruction tuning based and prompt engineering based approaches. Instruction-tuning approaches usually need to collect large-scale datasets and then further fine-tuned LLMs with parameter-efficient fine-tuning methods, such as LORA [26] and LongLORA [27]. TableLLAMA [28] and TAT-LLM [20] are typical examples of applying instruction tuning for the table processing. TableLLAMA is a generalist model for tables fine-tuned on Llama2 [29] with LongLORA. For fine-tuning TableLLAMA, a dataset collection named TableInstruct is proposed by collecting table-based samples from 14 datasets for 11 tasks. TAT-LLM is another fine-tuned LLAMA2 model specifically for the discrete reasoning over tables, which decomposes the Table-QA into three steps: Extractor, Reasoner and Executor. To construct the dataset for the model fine-tuning, TAT-LLM proposes a template to guide LLM in generating data following the proposed step-wise pipeline.
@@ III System Model
As discussed in Section I, in this study, we consider the scenarios in which IoT devices cannot access commercial LLMs because of data privacy and security considerations, which is a practical setting in many institutions, such as hospitals. Alternatively, as shown in Figure 1, an open-source LLM can be deployed in a local edge server to process the requests from IoT devices. Because fine-tuning an LLM for a specific task requires substantial datasets and computational resources, limiting the scalability of extending to multiple tasks, limiting the scalability of extending to multiple tasks, we propose to use prompting methods to enhance the capacities for different tasks with a Prompt Management Module, a Post-processing Module and a Task-specific Prompts Database, as shown in Figure 1. For a text-based request from IoT devices, such as a translation task to a sentence, the proposed Prompt Management Module is responsible for parsing the request and searching the task-specific prompt template and demonstrations regarding the task, as steps 1, 2 and 3 in Figure 1. After obtaining the task-specific prompt template, the Prompt Management Module constructs the final prompt and sends the request to the LLM. In our design, the capacities to deal with new tasks can be easily extended to the system by adding new prompt templates and demonstrations in the Task-specific Prompt Database without training or fine-tuning to the LLM. Since the results generated by the LLM are sometimes not ideal, we propose a Post-processing Module to process the results further. For example, some methods must prompt the LLM model multiple times to obtain the final results, whose intermediate results and Optional Task Request should be processed by the Post-processing Module. Besides, programming language-aided prompting methods, such as PAL [12] and PoT [13], generate the Python code instead of the final results, which means that the Post-processing Module should also be responsible for executing the generated Python code to obtain the final results. It is worth mentioning that, in the proposed solution, the IoT devices and the edge server can be easily connected with Wi-Fi and other wireless network protocols, and the communication between them can be easily implemented with HTTP protocol.
@@
Following these discussed steps, Figure 2 shows the detailed components of the proposed two modules and the detailed workflow of the proposed GIoT system. Specifically, the Prompt Management Module consists of three components: Request Parsing, Prompt Search, and Prompt Generation. The Request Parsing component receives and parses the requests from the IoT devices and outputs Task ID, Task Step and parsed Data, in which Task ID and Task Step would be used in the Prompt Search component to search corresponding Prompt Instructions and Prompt Demonstrations from the Task Specific Prompts Database. The parsed Data generated by the Request Parsing component is the information that needs to be further processed by the LLM. For example, for a service of translating English to Chinese, the request from the IoT devices can be ”Task Name: Translation. Data: This is a sample translation service.” Then the Data parsed by the Request Parsing component should be ”This is a sample translation service.”, which would be used by the Prompt Generation component to generate the final Task Prompt together with the Prompt Instruction and Prompt Demonstrations. It is worth mentioning that Prompt Generation can implement customize demonstration selection methods to optimize the performance of applying In Context Learning (ICL) [36]. For the Post-processing Module, since the output of the LLM cannot always follow the instructions, a Result Parsing function is needed to refine the outputs. For single-stage prompting methods, the result generated by the Result Parsing function can be the final result returned to the IoT device, as the path 1 shown in Figure 2. By contrast, prompting methods, such as PoT [13] and PAL [12], generate Python code and conduct the reasoning steps by an external Python Interpreter. Therefore, a Code Execution component should be included in the Post-processing Module. At last, a Optional Request Management component need to be implemented for the prompting methods with multiple-stages.
@@ IV-A Problem Formulation and Analysis
For a Table-QA service deployed in the proposed LLM-based GIoT system, it contains two key components: a parametric a LLM with parameters and a Retriever , in which the Retriever is responsible for selecting proper demonstrations used in ICL. Therefore, for a input query pair, , including a query table and a query question, the output answer can be .
@@ IV-B2 Demonstration Crafting and Selection
As shown in Figure 5, three prompts need to be constructed to answer a question; the demonstrations used in these three prompts are critical for the LLM’s performance, especially for the task-conducting stage to generate the reasoning function. Therefore, we defined a series of atomic operations to describe the reasoning logic to answer questions, as shown in Table I. For the task-planning step, we craft a question-and-answer pair for each type of operation and instruct the LLM to generate the Relevant Columns, Operations and Programming Steps, as shown in Figure 6. For the task-conducting prompts, we construct two question and reasoning function pairs for each defined operation and use operation as the metric to select these pairs. Even though we prompt the LLM to select one defined operation as output, the LLM can generate results without following the instructions. Therefore, when the Operations cannot be matched, the default setting contains all the question and function pairs. Figure 7 shows an example of COUNT operation, which includes a Meta Information table, Column Details and two questions with their Python solutions and default answers. It is worth mentioning that the demonstrations and prompts discussed in this section are stored in the Task-specific Prompts Database, and using operations as metrics to select proper demonstrations is the function of the proposed Prompt Management Module, as shown in Figures 1 and 2.
@@ V-A Datasets and Experimental Settings
We evaluate our proposed prompting solution on WikiTableQA [47] and TabFact [55] datasets, which are two datasets created by Wiki-tables without text context. WikiTableQA mainly contains compositional questions, such as questions requiring counting and ranking table contents. TabFact is a Fact Verification dataset, which can be treated as a special setting of a typical Table-QA problem whose answer set is . Since the proposed solution of this study is based on ICL, which is a few-shot learning setting without any training stage, we only use the test set of these two datasets to evaluate the performance, which contains 4344 and 12828 QA pairs, respectively. TabFact dataset further categorizes the test set into simple and complex groups, which include 4219 and 8609 QA pairs, respectively. Considering the large size of TabFact dataset, some studies [31, 35] conducted experiments one a small subset of TabFact, which contains 1,005 simple and 1,019 complex QA pairs. To compare with these studies, we also report the results on this small test subset of TabFact.
@@
As the proposed LLM-based GIoT system is designed to be deployed in a local network, we use open-source LLMs, including Mixtral-8x7B 111https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, Mistral-7B 222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, DeepSeek-67B 333https://huggingface.co/deepseek-ai/deepseek-llm-67b-chat and DeepSeek-7B 444https://huggingface.co/deepseek-ai/deepseek-llm-7b-chat to conduct our experiments. Since the proposed solution in this study is a prompt engineering method, we include Direct Prompting, CoT [14], PoT [13], Binder [35] and Dater [31] as benchmarks, in which Binder and Dater are two solutions leveraging SQL. For the implementation of benchmark methods, we use the implementation of TableCoT555https://github.com/wenhuchen/TableCoT [30] for the Direct Prompting and CoT. We re-implemented the PoT method following the example prompts reported in PoT [13]. The results of Dater and Binder are directly from study [31]. We use beam search decoding for the experiments. At last, even though we employ ICL to provide demonstrations to guide the LLM output of the final results in a ”{},” sometimes LLMs can fail to follow this output format. Therefore, we use a simple answer alignment step to post-process the results with incorrect formats. More specifically, we employ a direct prompting method by providing a few demonstrations containing the question, answer and formatted answer following TableCoT [30]. It is worth mentioning that we use the default precision of parameters, namely bfloat16, for the LLMs in this section.
@@ V-B1 Ablation Study
As discussed in Section IV, our proposed prompting solution consists of three stages: task-planning, task-conducting and task-correction. The task-planning stage can output the relevant columns and reasoning steps to answer the query question, which can be treated as a step of table decomposition and question decomposition, which can also be applied to the conventional PoT. In the task-conducting stage, we prompt the LLM to generate the Python code and a default answer. The default answer is the final answer when the Python code fails to run even after the task-correction stage. Since the default answer is generated after the Python code, it can be treated as an implicit CoT where the Python code is the reasoning rationales in the CoT. Finally, the third stage generates normalization functions to correct the errors in the Python code caused by the heterogeneous data types, which rely on the relevant columns generated by the first stage. Therefore, in this section, we conduct four ablation experiments by applying task-planning, default answer in task-conducting, task-planning and task-correction, and task-planning and default answer in task-conducting to the conventional PoT. The experimental results are shown in Table IV, where Plan, Correction and Default represent applying task-planning, task-correction and the default answer in task-conducting stages. We use Mixtral-8x7B as the LLM, and the experimental results demonstrate that each of the proposed three components can significantly improve the PoT baseline.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capacities on various tasks, and integrating the capacities of LLMs into the Internet of Things (IoT) applications has drawn much research attention recently. Due to security concerns, many institutions avoid accessing state-of-the-art commercial LLM services, requiring the deployment and utilization of open-source LLMs in a local network setting. However, open-source LLMs usually have more limitations regarding their performance, such as their arithmetic calculation and reasoning capacities, and practical systems of applying LLMs to IoT have yet to be well-explored. Therefore, we propose an LLM-based Generative IoT (GIoT) system deployed in the local network setting in this study. To alleviate the limitations of LLMs and provide service with competitive performance, we apply prompt engineering methods to enhance the capacities of the open-source LLMs, design a Prompt Management Module and a Post-processing Module to manage the tailored prompts for different tasks and process the results generated by the LLMs. To demonstrate the effectiveness of the proposed system, we discuss a challenging Table Question Answering (Table-QA) task as a case study of the proposed system, as tabular data is usually more challenging than plain text because of their complex structures, heterogeneous data types and sometimes huge sizes. We conduct comprehensive experiments on two popular Table-QA datasets, and the results show that our proposal can achieve competitive performance compared with state-of-the-art LLMs, demonstrating that the proposed LLM-based GIoT system can provide competitive performance with tailored prompting methods and is easily extensible to new tasks without training.
@@
Despite the remarkable capacities of LLMs, many issues need to be considered when integrating LLMs into IoT systems for an LLM-based GIoT system. First, following the Scaling Law [9], the capacities of an LLM are growing with its number of parameters and the scale of the training dataset, making LLMs often have a tremendous number of parameters, which leads to high hardware requirements for the training and inference. For example, popular open-source LLMs providing competitive performance on some public benchmarks often have around 70 billion parameters, such as Mixtral-8x7B [10] and Llama-3-70B [11]. These numbers of parameters make it not practical to deploy LLMs on edge IoT devices. Even though deploying LLMs on the edge and cloud servers can be an option, the efficiency of an LLM-based GIoT system still needs to be carefully considered. Besides the hardware requirements and efficiency issues, data privacy and security issues hinder many institutions from accessing commercial state-of-the-art LLMs [12, 13]. For example, for a healthcare IoT application collecting private data from patients, it is necessary to avoid uploading collected data to public, commercial LLM services for data analysis because of privacy issues. Therefore, a safe, transparent, and controllable open-source LLM deployed in a local network is more suitable for many institutions, even though commercial LLMs can provide better services. At last, the performance and the scalability of an LLM-based GIoT system are another two critical considerations because both commercial and open-source LLMs have some inherent limitations [14, 15, 16], such as their hallucination issues, limited reasoning capacities for complex tasks. Typically, prompting methods and fine-tuning are two directions that can alleviate these limitations. Specifically, prompting methods improve the performance of an LLM by designing tailored prompts for different tasks to elicit the capacities of an LLM. For example, Chain of Thought (CoT) [16] is a popular prompting method to improve the LLM’s reasoning performance by providing reasoning rationales. Program of Thoughts (PoT) [15] is another prompting method generating Python code to offload the reasoning and calculation tasks to the Python interpreter. Since these prompting methods focus on tailoring task-specific prompts for LLMs, they often lead to longer inference time because of their larger number of prompting tokens. By contrast, fine-tuning an LLM for a task can also improve the performance but requires labeled datasets and computation resources for the model training. As discussed in some studies [17, 18], fine-tuning an LLM for unseen tasks can increase the model’s bias and degrade its general capacities.
@@ II-B Table Question Answering
Since this study uses the Table-QA task as the case study of the proposed LLM-based GIoT system, we include recent studies using LLM to solve the Table-QA problem in this section. Specifically, most of the studies applying LLMs to the Table-QA task can be categorized into instruction tuning based and prompt engineering based approaches. Instruction tuning approaches usually need to collect large-scale datasets and then further fine-tuned LLMs with parameter-efficient fine-tuning methods, such as LORA [28] and LongLORA [29]. TableLLAMA [30] and TAT-LLM [22] are typical examples of applying instruction tuning for the table processing. TableLLAMA is a generalist model for tables fine-tuned on Llama2 [31] with LongLORA. For fine-tuning TableLLAMA, a dataset collection named TableInstruct is proposed by collecting table-based samples from 14 datasets for 11 tasks. TAT-LLM is another fine-tuned LLAMA2 model specifically for the discrete reasoning over tables, which decomposes the Table-QA into three steps: Extractor, Reasoner and Executor. To construct the dataset for the model fine-tuning, TAT-LLM proposes a template to guide LLM in generating data following the proposed step-wise pipeline.
@@ III System Model
As discussed in Section I, in this study, we consider the scenarios in which IoT devices cannot access commercial LLMs because of data privacy and security considerations, which is a practical setting in many institutions, such as hospitals. Alternatively, as shown in Figure 1, an open-source LLM can be deployed in a local edge server to process the requests from IoT devices. Because fine-tuning an LLM for a specific task requires substantial datasets and computational resources, limiting the scalability of extending to multiple tasks, limiting the scalability of extending to multiple tasks, we propose to use prompting methods to enhance the capacities for different tasks with a Prompt Management Module, a Post-processing Module and a Task-specific Prompts Database, as shown in Figure 1. Typically, a prompt often consists of instructions and demonstrations, in which the instructions are used to describe the task, and the demonstrations are the examples provided to the LLMs to show the desired outputs. For a text-based request from IoT devices, such as a translation task to a sentence, the proposed Prompt Management Module is responsible for parsing the request and searching for the task-specific prompt instructions and demonstrations regarding the task, as steps 1, 2 and 3 in Figure 1. After obtaining the task-specific prompt instructions and demonstrations, the Prompt Management Module constructs the final prompt and sends the request to the LLM. In our design, the capacities to deal with new tasks can be easily extended to the system by adding new prompt instructions and demonstrations in the Task-specific Prompt Database without training or fine-tuning to the LLM [46]. Since the results generated by the LLM are sometimes not ideal, we propose a Post-processing Module to process the results further. For example, some methods must prompt the LLM model multiple times to obtain the final results, whose intermediate results and Optional Task Request should be processed by the Post-processing Module. Besides, programming language-aided prompting methods, such as PAL [14] and PoT [15], generate the Python code instead of the final results, which means that the Post-processing Module should also be responsible for executing the generated Python code to obtain the final results. It is worth mentioning that, in the proposed solution, the IoT devices and the edge server can be easily connected with Wi-Fi and other wireless network protocols, and the communication between them can be easily implemented with HTTP protocol.
@@
Following these discussed steps, Figure 2 shows the detailed components of the proposed two modules and the detailed workflow of the proposed GIoT system. Specifically, the Prompt Management Module consists of three components: Request Parsing, Prompt Search, and Prompt Generation. The Request Parsing component receives and parses the requests from the IoT devices and outputs Task ID, Task Step and parsed Data, in which Task ID and Task Step would be used in the Prompt Search component to search corresponding Prompt Instructions and Prompt Demonstrations from the Task-specific Prompts Database. The parsed Data generated by the Request Parsing component is the information that needs to be further processed by the LLM. For example, for a service of translating English to Chinese, the request from the IoT devices can be ”Task Name: Translation. Data: This is a sample translation service.” Then the Data parsed by the Request Parsing component should be ”This is a sample translation service.”, which would be used by the Prompt Generation component to generate the final Task Prompt together with the Prompt Instruction and Prompt Demonstrations. It is worth mentioning that Prompt Generation can implement customize demonstration selection methods to optimize the performance of applying In Context Learning (ICL) [38]. For the Post-processing Module, since the output of the LLM cannot always follow the instructions, a Result Parsing function is needed to refine the outputs. For single-stage prompting methods, the result generated by the Result Parsing function can be the final result returned to the IoT device, as the path 1 shown in Figure 2. By contrast, prompting methods, such as PoT [15] and PAL [14], generate Python code and conduct the reasoning steps by an external Python Interpreter. Therefore, a Code Execution component should be included in the Post-processing Module. At last, an Optional Request Management component need to be implemented for the prompting methods with multiple-stages.
@@ IV-A Problem Formulation and Analysis
A Table-QA service deployed in the proposed LLM-based GIoT system contains two key components: a parametric LLM with parameters and a Retriever , in which the Retriever is responsible for selecting proper instructions and demonstrations used in ICL. Therefore, for an input query pair, , including a query table and a query question, the output answer after rounds of prompting can be defined as:
where is an empty string.
@@ IV-B2 Demonstration Crafting and Selection
As shown in Figure 5, three prompts need to be constructed to answer a question; the demonstrations used in these three prompts are critical for the LLM’s performance, especially for the task-conducting stage to generate the reasoning function. Therefore, we defined a series of atomic operations to describe the reasoning logic to answer questions, as shown in Table I. For the task-planning step, we craft a question-and-answer pair for each type of operation and instruct the LLM to generate the Relevant Columns, Operations and Programming Steps, as shown in Figure 6. For the task-conducting prompts, we construct two question and reasoning function pairs for each defined operation and use operation as the metric to select these pairs. Even though we prompt the LLM to select one defined operation as output, the LLM can generate results without following the instructions. Therefore, when the Operations cannot be matched, the default setting contains all the question and function pairs. Figure 7 shows an example of COUNT operation, which includes a Statistics Table, Column Details and two questions with their Python solutions and default answers. It is worth mentioning that the demonstrations and prompts discussed in this section are stored in the Task-specific Prompts Database, and using operations as metrics to select proper demonstrations is the function of the proposed Prompt Management Module, as shown in Figures 1 and 2.
@@ V-A Datasets and Experimental Settings
We evaluate our proposed prompting solution on WikiTableQA [50] and TabFact [58] datasets, which are two datasets created by Wiki-tables without text context. WikiTableQA mainly contains compositional questions, such as questions requiring counting and ranking table contents. TabFact is a Fact Verification dataset, which can be treated as a special setting of a typical Table-QA problem whose answer set is . Since the proposed solution of this study is based on ICL, which is a few-shot learning setting without any training stage, we only use the test set of these two datasets to evaluate the performance, which contains 4344 and 12828 QA pairs, respectively. TabFact dataset further categorizes the test set into simple and complex sets, which include 4219 and 8609 QA pairs, respectively. For the simple set, the QA pairs are usually obtained from a single row or record in the table, reflecting unary facts without complex logical inference. By contrast, for the complex set, the QA pairs are created by information from multiple columns and rows and derived by complex semantic operations, such as argmax, argmin, and the table records are also rewritten to include more semantic understanding. Considering the large size of TabFact dataset, some studies [33, 37] conducted experiments one a small subset of TabFact, which contains 1,005 simple and 1,019 complex QA pairs. To compare with these studies, we also report the results on this small test subset of TabFact.
@@
As the proposed LLM-based GIoT system is designed to be deployed in a local network, we use open-source LLMs, including Mixtral-8x7B 111https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, Mistral-7B 222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, DeepSeek-67B 333https://huggingface.co/deepseek-ai/deepseek-llm-67b-chat and DeepSeek-7B 444https://huggingface.co/deepseek-ai/deepseek-llm-7b-chat to conduct our experiments. Specifically, we deploy the LLMs with vLLM [59] on a workstation with 8 NVIDIA A100 40G GPUs. For the Mistral-7B and DeepSeek-7B models with default 16-bit parameters, one A100 40G GPU has provided enough RAM. By contrast, for the Mixtral8x7B and DeepSeek-67B with default 16-bit parameters, four A100 40G GPUs are needed for their deployment with tensor parallel.
Since the proposed solution in this study is a prompt engineering method, we include Direct Prompting, CoT [16], PoT [15], Binder [37] and Dater [33] as benchmarks, in which Binder and Dater are two solutions leveraging SQL. For the implementation of benchmark methods, we use the implementation of TableCoT555https://github.com/wenhuchen/TableCoT [32] for the Direct Prompting and CoT. We re-implemented the PoT method following the example prompts reported in PoT [15]. The results of Dater and Binder are directly from study [33]. We use a simple beam search decoding for the experiments. It is worth mentioning that self-consistency and more complex sampling decoding methods are often used in state-of-the-art techniques, such as Dater and Binder, to improve the quality of generated texts, which can introduce significant inference overhead. At last, even though we employ ICL to provide demonstrations to guide the LLM output of the final results in a ”{},” sometimes LLMs can fail to follow this output format. Therefore, we use a simple answer alignment step to post-process the results with incorrect formats. More specifically, we employ a direct prompting method by providing a few demonstrations containing the question, answer and formatted answer following TableCoT [32]. It is worth mentioning that we use the default precision of parameters, namely bfloat16, for the LLMs in this section.
@@ V-B1 Ablation Study
As discussed in Section IV, our proposed prompting solution consists of three stages: task-planning, task-conducting and task-correction. The task-planning stage can output the relevant columns and reasoning steps to answer the query question, which is also a step of table decomposition and question decomposition that can be applied to the conventional PoT. In the task-conducting stage, we prompt the LLM to generate the Python code and a default answer. The default answer is the final answer when the Python code fails to run even after the task-correction stage. Since the default answer is generated after the Python code, it can be treated as an implicit CoT where the Python code is the reasoning rationales in the CoT. Finally, the third stage generates normalization functions to correct the errors in the Python code caused by the heterogeneous data types, which rely on the relevant columns generated by the first stage. Therefore, in this section, we conduct four ablation experiments by applying task-planning, default answer in task-conducting, task-planning and task-correction, and task-planning and default answer in task-conducting to the conventional PoT. The experimental results are shown in Table IV, where Plan, Correction and Default represent applying task-planning, task-correction and the default answer in task-conducting stages. We use Mixtral-8x7B as the LLM, and the experimental results demonstrate that each of the proposed three components can significantly improve the PoT baseline.